This year, Sooner Company reports current E&P of negative $300,000. Its accumulated E&P at the beginning of the year was $200,000. Sooner distributed $400,000 to its sole shareholder, Boomer Wells, on June 30 of this year. Boomer’s tax basis in his Sooner stock is $75,000.

a. How much of the $400,000 distribution is treated as a dividend to Boomer?

b. What is Boomer’s tax basis in his Sooner stock after the distribution?

c. What is Sooner’s balance in accumulated E&P on the first day of the next year?
Answer #1

Part A

Deficit in current E&P = 300000*365*181= 148767

Thus, Boomer reports dividend = 200000-148767 = $51233

Part B

Boomer's tax basis in Sooner stock = $0

Tax basis = beginning basis in stock (75000) - lesser of distribution in excess of stock basis (348767) or basis in Sooner's stock (75000).

Part C

Sooner's balance in accumulated E&P = 300000/365*183 = $150411
Answer #2

a) $50,000 

Defecit in Current E&P = (current E&P/Number of months in a year) * ( number of months ( June 30-dec 31)
a1) ($300,000/12)*(6)= 150,000
a2) Dividend to MR. BM= Accumulated E&P- Deficit in Current E&P= $200,000-$150,000= $50,000.

b) 0
i). Accumulated E&P= (400,000-(300,000-200,000)= 300,000
ii) Tax basis= 75,000< 300,000
75,000-75,000=0
The remaining (300,000-75,000)= 275,000 is treats as a capital gain

c)(150,000)
(accumulated E&P + Current E&P)- Dividends Paid= 200,000+(300,000)-50,000=(150,000)